Prove that the line x − 2y + 4a = 0 touches the parabola y2= 4ax, and find the coordinates of P, the point of contact. If the line x − 2y + 2a = 0 meets the parabola in Q, R, and M is the mid-point of QR, prove that PM is parallel to the axis of x, and that this axis and the line through M perpendicular to it meet on the normal at P to the parabola.


1
Expert's answer
2021-08-12T17:53:24-0400
"x \u2212 2y + 4a = 0"

"y^2= 4ax"

Then


"y^2=4a(2y-4a)"

"y^2-8ay+16a^2=0"

"(y-4a)^2=0"

"y=4a"

Since the equation "y^2=4a(2y-4a)" has the only root, then the line "x \u2212 2y + 4a = 0" touches the parabola "y^2=4ax."


"y=4a, x=(4a)^2\/4a=4a"

The point of contact "P(4a, 4a)."


Suppose that "Q" has coordinates "(aq^2, 2aq)" and "R" has coordinates "(ar^2, 2ar)."

The line "x \u2212 2y + 2a = 0" meets the parabola in "Q, R"


"aq^2-4aq+2a=0"

"q^2-4q+2=0, a\not=0"

"q=2\pm\sqrt{2}"


"ar^2-4ar+2a=0"

"r^2-4r+2=0, a\not=0"

"r=2\pm\sqrt{2}"




"(2-\sqrt{2})^2=6-4\sqrt{2}"

"(2+\sqrt{2})^2=6+4\sqrt{2}"

Suppose that "Q" has coordinates "(a(6-4\sqrt{2}), a(4-2\sqrt{2}))" and "R" has coordinates "(a(6+4\sqrt{2}), a(4+2\sqrt{2}))."

Since "M" is the mid-point of "QR," "M" has coordinates "( 6a, 4a)."

Then the equation of "PM" is "y=4a."

Hence "PM" is parallel to the axis of "x."


Find the equation of the normal at "P" to the parabola


"slope=\dfrac{-1}{\dfrac{1}{2}}=-2"


"y=-2x+b"

"4a=-2(4a)+b=>b=12a"


"y=-2x+12a"

The lines "y=0, x=6a," and "y=-2x+12a" meet at the point "(6a, 0)."